The vapor pressure of liquid lead is 400. mm Hg at 1900 K. Assuming that its molar heat of vaporization is constant at 179 kJ/mol, the vapor pressure of liquid Pb is
mm Hg at a temperature of 1860 K.
1
Expert's answer
2019-11-11T08:29:00-0500

Solution.

"\Delta H(vap.) = \frac{R \times T1 \times T2 \times ln(\frac{p2}{p1})}{T2-T1}"

"p2 = p1 \times \exp(\frac{\Delta H(vap.) \times (T2-T1)}{R \times T1 \times T2}"

"p2 = 400 \times \exp(\frac{149*10^3 \times (1860-1900)}{8.31 \times 1900 \times 1860}"

p2 = 313.46 mm Hg

Answer:

p2 = 313.46 mm Hg